OpenStudy (deadshot):

What is the remainder when (3x^4 + 2x^3 - x^2 + 2x - 9) / (x + 2) ?

OpenStudy (dumbcow):

remainder theorem says if dividing f(x) by (x-c) then remainder is f(c) x-c= x+2 c = -2 what is f(-2)

OpenStudy (deadshot):

15?

OpenStudy (dumbcow):

yep, thats the remainder

OpenStudy (deadshot):

Thanks!